# Teaching Large Language Models to Maintain Contextual Faithfulness via Synthetic Tasks and Reinforcement Learning

### v1

@@ 4 Experiments
$\Delta$ Compared to Vanilla. +24.3 +22.6 +71.3 +25.6 +28.9 +28.2 +22.6 +9.8 +6.9 +10.2 +7.3 +6.4 +22.6 +16.2
@@
$\Delta$ Compared to Vanilla. +15.1 +14.2 +64.9 +15.0 +11.9 +8.2 +14.4 +4.4 +4.9 +9.0 +28.3 +4.0 +19.0 +12.3
[…]
@@
Table 1: Experimental results (%) on eleven datasets. The FollowRAG results represent the results averaged over these four open-domain QA datasets as shown in Table 7, including NaturalQA, TriviaQA, HotpotQA, and WebQSP. Bold numbers indicate the best performance of models with the same model size. Avg EM/Acc represents the average score between short-form task metrics (EM/Acc) and long-form task metric FaithScore (FS).
@@ 4.1 Tasks and Datasets
Short-form Generation Tasks. For short-form generation tasks, we use two counterfactual QA datasets (ConFiQA (Bi et al., 2024) and CNQ (Longpre et al., 2021)), a multiple-choice questions dataset FaithEval (Ming et al., 2025), and a factual QA dataset FiQA (Bi et al., 2024) that is the factual version of ConFiQA. These datasets ensure the answers appear in the contexts to evaluate the faithfulness. We also evaluate our method on four open-domain QA datasets within the FollowRAG benchmark (Dong et al., 2024) to evaluate the abilities of LLMs in real-world RAG scenarios, including NaturalQA (Kwiatkowski et al., 2019b), TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and WebQSP (Yih et al., 2016). In real-world RAG scenarios, the answer may not appear in the retrieved passages, and these passages tend to be noisy. We evaluate models based on whether gold answers are included in the generated responses (i.e., Acc) following Asai et al. (2024) and exact matching (EM) for QA tasks. For multiple-choice questions, we follow Ming et al. (2025) and use keyword matching to verify the accuracy.
Long-form Generation Tasks. We include a text summarization task XSum (Narayan et al., 2018), a text simplification task WikiLarge (Zhang and Lapata, 2017), and a long-form QA task CLAPNQ (Rosenthal et al., 2025). To evaluate the faithfulness of generated long-form answers, called FaithScore (FS), we use MiniCheck (Tang et al., 2024) to check whether the model response is grounded in the provided context. MiniCheck is a state-of-the-art method to recognize if LLM output can be grounded in given contexts. If the model response contains at least one statement that cannot be inferred from the context, we consider it as a negative response; otherwise, it is a positive response. We also query GPT-4o to evaluate the quality of generated responses, namely QualityScore.
More details are available in the Appendix C.
@@ 4.2 Baselines and Implementation Details
Baselines. We compare several baselines, including (1) Vanilla LLMs: including LLaMA-3-Instruct (Grattafiori et al., 2024) and Qwen-2.5-Instruct (Yang et al., 2024) of different sizes. We also conduct supervised fine-tuning on synthesized 10,000 short-form data as SFT baselines; (2) SOTA LLMs: We further evaluate the most advanced LLMs, including GPT-4o, GPT-4o-mini, OpenAI o1 (Jaech et al., 2024), Claude 3.7 Sonnet (Anthropic, 2025), Claude 3.7 Sonnet-Thinking, Deepseek R1, and Deepseek V3 (DeepSeek-AI et al., 2025a, b); (3) The Designed Methods to Improve Faithfulness of LLMs: Context-DPO (Bi et al., 2024) aligns LLMs through DPO with constructed faithful and unfaithful short-form answers, thus improving the faithfulness in short-form generation. SCOPE (Duong et al., 2025) introduces a pipeline to generate self-supervised task-specific data and applies preference training to enhance the faithfulness in a special task. We train it on the sampled training set of the summarization task XSum as SCOPEsum, regarding it as the method designed to improve the faithfulness of long-form response generation.
Implementation Details. Our main experiments are conducted on LLaMA-3-Instruct and Qwen-2.5-Instruct. More implementation details are shown in Appendix D, e.g., hyperparameters.
@@ 4.3 Main Results
Canoe Improves the Faithfulness of LLMs in Both Short-form and Long-form Response Generation. As shown in Table 1, Canoe shows consistent and significant improvements on 11 datasets measuring faithfulness. Canoe achieves substantial improvements in the overall score compared to original LLMs, e.g., 22.6% for Llama3-8B and 19.0% for Qwen2.5-7B in Avg EM score. Canoe also surpasses the most advanced LLMs (e.g., GPT-4o) in the overall score (both Avg EM and Avg Acc scores). This shows that Canoe can effectively align LLMs to be context-faithful. Meanwhile, for real-world RAG scenarios, our proposed Canoe can also improve the performance even though the answer may not appear in the retrieved passages, and these passages are often noisy.
@@
Table 2: QualityScore on long-form generation tasks.
@@
Canoe Maintains the Factuality of LLMs. We further evaluate whether Canoe will reduce the factuality of LLMs. Following Ming et al. (2025), we modify the original FaithEval and make it a closed-book QA setting, where no context is provided and LLMs need to give factual answers. In this case, the models rely entirely on their parametric knowledge of common facts, and we find that our proposed Canoe maintains the factuality compared to the untuned LLM as shown in Figure 3. However, when a new context with counterfactual evidence that contradicts the model’s parametric knowledge is introduced, performance declines sharply. For example, GPT-4o achieves 96.3% accuracy on factual closed-book QA task but only 47.5% on counterfactual QA task that evaluates the faithfulness of LLMs. This highlights that, unlike factuality, the faithfulness of LLMs is difficult to improve by simply scaling model parameters, which further indicates the necessity of a post-training method to improve faithfulness.
Canoe Improves the Quality of Long-form Response Generation. As shown in Table 2, we can find that our proposed Canoe also improves the quality of generations. This is because the proxy reward implicitly requires LLMs to generate easy-to-understand responses, which further optimizes the response quality. Canoe consistently improves the generation quality in the three long-form tasks, which illustrates the effectiveness of our method.
@@
[…]
@@
Table 4: Results of ablation study. EM/Acc in short-form tasks represents the average score between QA metrics (EM/Acc) and the accuracy of FaithEval. N/A means a false generation pattern hacks this metric.
@@ 4.4 Analysis
Ablation Study. We conduct an ablation study in Table 4. The result reveals that our proposed Canoe (including Dual-GRPO and the designed short-form data synthesis) significantly improves the faithfulness of LLMs in both short-form and long-form generation. For Dual-GRPO, we observe that directly applying GRPO to synthesized short-form data leads to over-optimizing short-form generation and a false response generation pattern. We find that tuned models tend to directly copy text spans from the given context as the final answer instead of following instructions in long-form generation tasks (we show the case study in the Appendix F.5 to visually show this phenomenon). Thus, the generated responses do not contain syntactically and semantically complete sentences for long-form generation tasks, which leads to low QualityScore performance and also invalidates the metric used for evaluating faithfulness. We also explore the effectiveness of our training data construction strategy. For the designed QA tasks used to ensure the complexity and diversity of training data, we replace the designed Reasoning-required Context and Inconsistent Context samples with the same number of samples that contain Straightforward Context. We find that involving these more challenging instances can improve the effectiveness of RL training. We also replace the data points that contain Counterfactual Context with the same number of factual samples. The designed Counterfactual Context improves the final performance as it prevents models from depending on their learned factual knowledge to find the right answers.
@@
[…]
@@
Case Study. We further conduct a case study in Table 4.4 to visually show the advantages of Canoe. Our method ensures the statements are faithful and comprehensive, and the text flows naturally.
Human Evaluation. Evaluating long-form generation tasks remains challenging (Li et al., 2024). Thus, we conduct human evaluation in the Appendix E to show the effectiveness of our method.
Discussion. We also discuss some possible concerns about Canoe in the Appendix F, e.g., the effect of the amount of synthesized data.
@@ 5 Conclusion
Limitations
Although experiments have confirmed the effectiveness of the proposed Canoe, four major limitations remain. Firstly, Canoe synthesizes short-form QA data and uses the proposed Dual-GRPO to improve the faithfulness of LLMs in long-form response generation implicitly; thus, how to directly synthesize long-form data and improve the faithfulness remains under-explored. Meanwhile, the synthesized short-form QA data is single-turn; thus, exploring the synthesis of multi-turn QA data presents an attractive direction for future research. The motivation behind our work is to improve the faithfulness of LLMs without human annotation, but it is still worth exploring how to incorporate the existing manually labeled data to further improve the faithfulness of the model. Finally, while our method achieves strong results, exploring additional strategies, e.g., using cold-start to get a better initial policy model and improve the reward scores in training for better performance across different downstream tasks is also a promising direction.
@@
  • In Section A, we report the details of constructing training data, e.g., the used triples and introduction of four designed tasks.
  • In Section B, we go into detail about the proposed Dual-GRPO, including the system prompt and formal expressions of three different well-designed rewards.
  • In Section C, we show the details of evaluations, e.g., the introduction of the used benchmarks and evaluation prompts.
  • In Section D, we show the details of our implementation and training, e.g., hyperparameters and the used GPUs.
  • In Section E, we show the implementation details of human evaluation.
  • In Section F, we discuss some possible questions about the proposed Canoe. For example, we discuss the effect of the amount of synthesized short-form data for RL training.
@@ A.1 Triples from Wikidata
To ensure the usability of the synthetic data and collected triples, we follow Bi et al. (2024) to collect entities corresponding to the top 1,000 most-visited Wikipedia pages from 2016 to 2023 and 41 relations selected by Bi et al. (2024) shown in Table 12. The most-visited Wikipedia pages are based on monthly page views and retain the most popular entities using criteria such as the number of hyperlinks. We finally collected 6,316 entities and 30,762 triples. We randomly select these triples to synthesize our training data, and finally construct 10,000 samples as the final training data.
@@ A.2 Construction of Four Different Tasks
Straightforward Context. As shown in Sec. 3.1, we keep the original collected factual triple as input to query GPT-4o to synthesize the data $(c,q,a)$. The prompts for querying GPT-4o to obtain the generated questions and contexts can be found in Figure 7 and Figure 8. We finally keep 2,000 such samples in the synthesized 10,000 training data, i.e., 20% of the data.
Reasoning-required Context. We construct paths $[(h^{1},r^{1},t^{1}),...,(h^{n},r^{n},t^{n})]_{n\leq 4}$ from a sub-graph; more details can be found in Sec. 3.1. Then, we use the $n$-th tail entity $t^{n}$ as the ground truth answer and use the constructed paths to query GPT-4o to obtain the multi-hop context and question. The prompts for querying GPT-4o to obtain the generated questions and contexts can be found in Figure 9 and Figure 10. We finally keep 2,000 such samples in the synthesized 10,000 training data, i.e., 20% of the data.
@@
Counterfactual Context. A counterfactual context includes statements that go against common sense found in the collected triples. Specifically, we construct samples with counterfactual contexts below by modifying previously collected triples (of three types, including straightforward context, reasoning-required context, and inconsistent context). We replace the tail entity $t$ of the original collected triple with a similar but counterfactual entity $t^{cf}$, which is obtained by query GPT-4o using prompt “Generate me a noun for an entity that is similar to the {$t$} but different, and require the entity to exist in the real-world, please tell me the answer directly:”. Then, we query GPT-4o to generate questions and counterfactual contexts to construct counterfactual samples, using the counterfactual triples. The prompts used in constructing samples with counterfactual contexts are the same as the prompts used in constructing the three different tasks above. The reason we construct samples with counterfactual context in this way is that this prevents the model from learning the appropriate factual knowledge to answer the question correctly, rather than correctly exploiting the given contextual information. Therefore, we construct the same number of samples as the summed number of the three types above (including straightforward context, reasoning-required context, and inconsistent context), i.e., 5000 samples (50% of the data). Meanwhile, this task stresses the importance of keeping answers faithful in contexts, as it stops them from relying solely on the learned knowledge of LLMs to provide correct answers.
Type Num Avg Len
Straightforward Context. 2,000 186.3
Reasoning-required Context. 2,000 262.2
Inconsistent Context. 1,000 421.2
Counterfactual Context. 5,000 260.8
Table 6: Statistics of the training data. Num indicates the number of samples. Avg Len shows the average length of the samples, including the context and question.

### v3

@@ 3 Experiments
[2pt/3pt] $\Delta$ Compared to Vanilla. +24.3 +22.6 +71.3 +25.6 +28.9 +28.2 +22.6 +9.8 +6.9 +10.2 +7.3 +6.4 +22.6 +16.2
@@
[2pt/3pt] $\Delta$ Compared to Vanilla. +15.1 +14.2 +64.9 +15.0 +11.9 +8.2 +14.4 +4.4 +4.9 +9.0 +28.3 +4.0 +19.0 +12.3
[2pt/3pt] Canoe-Qwen-14B 85.7 87.4 87.8 88.5 81.8 84.2 67.4 46.1 54.6 75.7 91.1 68.4 75.5 77.2
@@
Table 1: Experimental results (%) on eleven datasets. The FollowRAG results represent the results averaged over these four open-domain QA datasets, as shown in the Appendix C, including NaturalQA, TriviaQA, HotpotQA, and WebQSP. Bold numbers indicate the best performance of models with the same model size. Avg EM/Acc represents the average score between short-form task metrics (EM/Acc) and long-form task metric FaithScore (FS).
@@ 3.1 Tasks and Datasets
Short-form Generation Tasks. For short-form generation tasks, we use two counterfactual QA datasets, including ConFiQA (bi2024contextdpoaligninglanguagemodels) and CNQ (longpre-etal-2021-entity), a multiple-choice questions dataset FaithEval (ming2025faitheval), and a factual QA dataset FiQA (bi2024contextdpoaligninglanguagemodels) that is the factual version of ConFiQA. These datasets ensure the answers appear in the contexts to evaluate the faithfulness. We also evaluate our method on four open-domain QA datasets within the FollowRAG benchmark (dong2024generalinstructionfollowingalignmentretrievalaugmented) to evaluate the abilities of LLMs in real-world retrieval-augmented generation (RAG) scenarios, including NaturalQA (47761), TriviaQA (joshi-etal-2017-triviaqa), HotpotQA (yang-etal-2018-hotpotqa), and WebQSP (Yih2016TheVO). In real-world RAG scenarios, the answer may not appear in the retrieved passages, and these passages tend to be noisy. We evaluate models based on whether gold answers are included in the generated responses (i.e., Acc) following asai2024selfrag and exact matching (EM) for QA tasks. For multiple-choice questions, we follow ming2025faitheval and use keyword matching to verify the accuracy.
Long-form Generation Tasks. We include a text summarization task XSum (narayan-etal-2018-dont), a text simplification task WikiLarge (zhang-lapata-2017-sentence), and a long-form QA task CLAPNQ (rosenthal-etal-2025-clapnq). To evaluate the faithfulness of generated long-form answers, called FaithScore (FS), we use MiniCheck (tang-etal-2024-minicheck) to check whether the model response is grounded in the provided context. MiniCheck is a state-of-the-art method to recognize if LLM output can be grounded in given contexts. If the model response contains at least one statement that cannot be inferred from the context, we consider it as a negative response; otherwise, it is a positive response. We also query GPT-4o to evaluate the quality of generated responses, namely QualityScore.
More details can be found in the Appendix C.
@@ 3.2 Baselines and Implementation Details
Baselines. We compare several baselines, including (1) Vanilla LLMs: including LLaMA-3-Instruct (llama3) and Qwen-2.5-Instruct (qwen2.5) of different sizes. We also conduct supervised fine-tuning on synthesized 10,000 short-form data as SFT baselines. (2) SOTA LLMs: We further evaluate the most advanced LLMs, including GPT-4o, GPT-4o-mini, OpenAI o1 (jaech2024openai), Claude 3.7 Sonnet (Claude3S), Claude 3.7 Sonnet-Thinking, DeepSeek R1, and DeepSeek V3 (deepseekai2025deepseekr1incentivizingreasoningcapability; deepseekai2025deepseekv3technicalreport). (3) The Methods to Improve Faithfulness: Context-DPO (bi2024contextdpoaligninglanguagemodels) aligns LLMs through DPO with constructed faithful and unfaithful short-form answers, thus improving the faithfulness in short-form generation. SCOPE (duong2025scope) proposes a pipeline to generate self-supervised task-specific data and applies preference training to enhance the faithfulness in a special task. We train it on the sampled training set of the summarization task XSum as SCOPEsum, regarding it as the method designed to improve the faithfulness of long-form response generation.
Implementation Details. Our experiments are conducted on LLaMA-3-Instruct and Qwen-2.5-Instruct. Details are shown in Appendix D, e.g., hyperparameters.
@@ 3.3 Main Results
Canoe Improves the Faithfulness of LLMs in Both Short-form and Long-form Response Generation. As shown in Table 1, Canoe shows significant improvements on 11 faithfulness-related benchmarks. Canoe achieves substantial improvements in the overall score compared to original LLMs, e.g., 22.6% for Llama3-8B and 19.0% for Qwen2.5-7B in Avg EM score. Canoe also surpasses the most advanced LLMs (e.g., GPT-4o) in the overall score (both Avg EM and Avg Acc scores). This shows that Canoe effectively aligns LLMs to be context-faithful. Also, for real-world RAG scenarios, our proposed Canoe can also improve the performance even though the answer may not appear in the retrieved passages, and these passages are often noisy.
@@
Table 2: QualityScore (%) on long-form generation tasks.
@@
Canoe Maintains the Factuality of LLMs. We further evaluate whether Canoe will reduce the factuality of LLMs. Following ming2025faitheval, we modify the original FaithEval and make it a closed-book QA setting, where no context is provided and LLMs need to give factual answers. In this case, the models rely entirely on their parametric knowledge of common facts, and we find that our proposed Canoe maintains the factuality compared to the untuned LLM as shown in Figure 3. However, when a new context with counterfactual evidence that contradicts the model’s parametric knowledge is introduced, performance declines sharply. For example, GPT-4o achieves 96.3% accuracy on factual closed-book QA task but only 47.5% on counterfactual QA task that evaluates the faithfulness of LLMs. This highlights that, unlike factuality, the faithfulness of LLMs is difficult to improve by simply scaling model parameters, which further indicates the necessity of a post-training method to improve faithfulness.
Canoe Improves the Quality of Long-form Response Generation. As shown in Table 2, we can find that our proposed Canoe consistently improves the generation quality in the three long-form tasks. This is because the proxy reward implicitly requires LLMs to generate easy-to-understand responses, which further optimizes the response quality.
@@
[2pt/3pt] -w/o. Dual-GRPO (i.e., original GRPO) 60.5 66.6 N/A 23.5
[2pt/3pt] -w/o. Reasoning-required Context. 63.7 69.4 71.7 75.3
@@
Table 4: Results (%) of ablation study. EM/Acc in short-form tasks represents the average score between QA metrics (EM/Acc) and the accuracy of FaithEval. N/A means a false generation pattern hacks this metric.
@@ 3.4 Analysis
Ablation Study. We conduct an ablation study in Table 4. The result reveals that our proposed Canoe (including Dual-GRPO and the designed short-form data synthesis) significantly improves the faithfulness of LLMs in both short-form and long-form generation tasks. For Dual-GRPO, we observe that directly applying GRPO to synthesized short-form data leads to over-optimizing short-form generation and a false response generation pattern. We find that tuned models tend to directly copy text spans from the given context as the final answer instead of following instructions in long-form generation tasks (we show the case study in Table 3.4 to visually show this phenomenon). Thus, the generated responses do not contain syntactically and semantically complete sentences for long-form generation tasks, which leads to low QualityScore performance and also invalidates the metric used for evaluating faithfulness. For the designed QA tasks used to ensure the complexity and diversity of training data, we replace the designed Reasoning-required Context and Inconsistent Context samples with the same number of samples that contain Straightforward Context. We find that involving these more challenging instances can improve the effectiveness of RL training. We also replace the data points that contain Counterfactual Context with the same number of factual samples. The designed Counterfactual Context improves the final performance as it prevents models from depending on their learned factual knowledge to find the right answers.
Case Study. We find that directly applying GRPO instead of our Dual-GRPO to synthesized short-form data leads to over-optimizing short-form generation and a false response generation pattern. As shown in Table 3.4, we find that the tuned model GRPO-LLaMA-8B tends to copy text spans from the given context as the final answer instead of following instructions in the summarization task. We also observe the same phenomenon in the tasks of text simplification and long-form QA; more details can be found in the Appendix F. However, when we apply Dual-GPRO, we find that trained models can generate fluent and complete sentences. Thus, Dual-GRPO not only improves the faithfulness of LLMs in two types of response generation but also ensures the utility. We also conduct a case study in Table 3.4 to show the advantages of Canoe. Canoe ensures the statements are faithful and comprehensive, and the text flows naturally.
Text Summarization Task
Inputs: {Documents}
Instruction: Refer to the passage above and provide a summary as the response.
[0.8pt/2pt] GRPO-LLaMA-8B: Comedy clubs. ()
[0.8pt/2pt] Canoe-LLaMA-8B: Kareem Badr and his friends were able to turn around the struggling comedy club, the Hideout, by applying their skills in problem-solving and efficiency. They broke down the business into small chunks and figured out how to make each aspect better and more efficient. This approach helped them to improve the club’s performance and eventually expand it. The passage also mentions that the growing popularity of stand-up comedy is fueled in part by younger audiences, who view humorous TV hosts as their primary source for news. Many comedy clubs now require talent, and there is a career and a living to be made by good comedians who are not superstars. The clubs also make money from selling drinks and dinners, as well as from cover charges and talent fees. In addition, digital media has given live comedy a boost, with one terrific joke being able to get a million hits online. ()
Table 5: Case study between the models trained via GRPO and our model for the summarization task.
@@
[0.8pt/2pt] Canoe-Qwen-7B: According to the information given, in the Three’s Company episode “Opening Night,” when Jack breaks the plates at the end of the celebration, he exclaims “Opa!”.() This suggests that “Opa!” might be a phrase commonly used in Greek culture during such festivities. () “Opa!” is often used to express joy, excitement, or approval, which aligns with the celebratory nature of plate smashing in Greek traditions. ()
@@
Human Evaluation. Evaluating long-form generation tasks remains challenging. Thus, we conduct human evaluation in the Appendix E to show the effectiveness. We find that our method reduces faithfulness hallucinations and also ensures the response quality for three long-form generation tasks.
Discussion. We discuss possible concerns in the Appendix F, e.g., the effect of the amount of synthesized data, multilingual transfer ability, and context length generalization of Canoe.
4 Related Work
Recently, the demand for utilizing LLMs to generate coherent text responses based on the provided contexts has continued to grow, particularly in text summarization and RAG scenarios. However, LLMs are often criticized for generating outputs that deviate from the provided contents, namely faithfulness hallucination (li2022faithfulnessnaturallanguagegeneration; ji-survey; hit-survey). Many approaches have been proposed to improve the faithfulness of language models (zhu-etal-2021-enhancing; jones2024teaching; hu2024mitigatinglargelanguagemodel; li2025hallucinationdilemmafactualityawarereinforcement; yang-etal-2025-longfaith; zhang-etal-2025-longreward; wang2025retrievalaugmentedgenerationconflictingevidence). The first line of work focuses on the inference stage, such as designing prompts to encourage context integration (zhou-etal-2023-context), improving context quality via explicit denoising (xu2024recomp), and context-aware decoding to amplify contexts (shi-etal-2024-trusting). Although effective, these approaches primarily serve as a compensatory way rather than enabling the model to inherently learn to prevent generating unfaithful responses. Thus, many studies attempt to apply post-training methods to improve the faithfulness. bi2024contextdpoaligninglanguagemodels utilizes constructed faithful and unfaithful short-form completions and applies DPO to align LLMs to be context-faithful in short-form QA tasks. huang2025improvingcontextualfaithfulnesslarge trains LLMs to discriminate between faithful and unfaithful responses in long-form QA tasks by unfaithful response synthesis and contrastive tuning. li2025generatediscriminateevolveenhancing introduces a self-evolving framework to maintain the faithfulness of LLMs in long-form QA tasks through iterative self-improvement. duong2025scope proposes a pipeline to generate a self-supervised task-specific dataset and applies preference training to enhance the faithfulness for a special task. However, these methods struggle to consistently improve the faithfulness of LLMs across various tasks, as these methods are designed for specific tasks. Thus, how to consistently improve the faithfulness of LLMs on different downstream tasks, including short-form and long-form generation tasks, still remains under-explored.
@@ 5 Conclusion
  • In Section A Training Data Details, we report the details of constructing training data, e.g., the used triples and introduction of four designed tasks.
  • In Section B Dual-GRPO Details, we go into detail about the proposed Dual-GRPO, including the system prompt and formal expressions of three well-designed rewards.
  • In Section C Evaluation Details, we show the details of evaluations, e.g., the introduction of the used benchmarks and evaluation prompts.
  • In Section D Implementations Details, we show the details of our implementation and training, e.g., hyperparameters.
  • In Section E Human Evaluation, we show the implementation details of human evaluation.
  • In Section F Discussion, we discuss some possible questions about the proposed Canoe. For example, we discuss the effect of the amount of synthesized data for training.
@@ A.1 Triples from Wikidata
To ensure the usability of the synthetic data and collected triples, we follow bi2024contextdpoaligninglanguagemodels to collect entities corresponding to the top 1,000 most-visited Wikipedia pages from 2016 to 2023 and 41 relations selected by bi2024contextdpoaligninglanguagemodels shown in Table 13. The most-visited Wikipedia pages are based on monthly page views and retain the most popular entities using criteria such as the number of hyperlinks. We finally collected 6,316 entities and 30,762 triples. We randomly select these triples to synthesize our training data, and finally construct 10,000 samples as the final training data.
@@ A.2 Construction of Four Different Tasks
Straightforward Context. As shown in Sec. 2.1, we keep the original collected factual triple as input to query GPT-4o to synthesize the data $(c,q,a)$. The prompts for querying GPT-4o to obtain the generated questions and contexts can be found in Figure 7 and Figure 8. We keep 2,000 such samples in the synthesized 10,000 training data, i.e., 20% of the data.
Reasoning-required Context. We construct paths $(h^{1},r^{1},t^{1}),...,(h^{n},r^{n},t^{n})]_{n\leq 4}$ from a sub-graph; more details can be found in Sec. 2.1. Then, we use the $n$-th tail entity $t^{n}$ as the ground truth answer and use the constructed paths to query GPT-4o to obtain the multi-hop context and question. The prompts for querying GPT-4o to obtain the generated questions and contexts can be found in Figure 9 and Figure 10. We finally keep 2,000 such samples in the synthesized 10,000 training data, i.e., 20% of the data.
@@
Counterfactual Context. A counterfactual context includes statements that go against common sense found in the collected triples. Specifically, we construct samples with counterfactual contexts below by modifying previously collected triples (of three types, including straightforward context, reasoning-required context, and inconsistent context). We replace the tail entity $t$ of the original collected triple with a similar but counterfactual entity $t^{cf}$, which is obtained by query GPT-4o using prompt “Generate me a noun for an entity that is similar to the {tt} but different, and require the entity to exist in the real-world, please tell me the answer directly:”. Then, we query GPT-4o to generate questions and counterfactual contexts to construct counterfactual samples, using the counterfactual triples. The prompts used in constructing samples with counterfactual contexts are the same as the prompts used in constructing the three different tasks above. The reason we construct samples with counterfactual context in this way is that this prevents the model from learning the appropriate factual knowledge to answer the question correctly, rather than correctly exploiting the given contextual information. Therefore, we construct the same number of samples as the summed number of the three types above (including straightforward context, reasoning-required context, and inconsistent context), i.e., 5000 samples (50% of the data). This task stresses the importance of keeping answers faithful in contexts, as it stops them from relying solely on the learned knowledge of LLMs to provide correct answers.
black Type Num Avg Len Straightforward Context. 2,000 186.3 Reasoning-required Context. 2,000 262.2 Inconsistent Context. 1,000 421.2 Counterfactual Context. 5,000 260.8 black
Table 7: Statistics of the training data. Num shows the number of samples. Avg Len shows the average length of the samples.